Find the equation of the circle. Touching the x-axis and passing through the points (3,1),(10,8). ty
Answer by Edwin McCravy(20086)

Find the equation of the circle. Touching the x-axis and passing through the points (3,1),(10,8).
We plot those points:




We can guess how to draw this circle using a compass.  There are
two possible circles:



Let the center of the circle be (h,k), then since the circle touches
the x-axis, we draw a green vertical radius down to the x-axis:



Because it touches the x-axis, we can see from the graph above, k,
which is the center's y-coordinate is the same as the radius r.

The equation of a crcle with center (h,k) and radius r has equation:

(x - h)² + (y - k)² = r²

And since k = r, it is

(x - h)² + (y - r)² = r²

x² - 2hx + h² + y² - 2ry + r² = r²

x² - 2hx + h² + y² - 2ry = 0 

We now substitute each of the given points in the equation:

Substituting (x,y) = (3,1)

3² - 2h(3) + h² + 1² - 2r(1) = 0

9 - 6h + h² + 1 - 2r = 0

10 - 6h + h² - 2r = 0

Substituting (x,y) = (10,8)

x² - 2hx + h² + y² - 2ry = 0 

10² - 2h(10) + h² + 8² - 2r(8) = 0

100 - 20h + h² + 64 - 16r = 0

164 - 20h + h² - 16r = 0




So we have this system of equations:

 10 -  6h + h² -  2r = 0
164 - 20h + h² - 16r = 0
 
We eliminate r by multiplying the first equation by -8
and adding the two equations:

-80 + 48h - 8h² + 16r = 0
164 - 20h +  h² - 16r = 0
-------------------------
 84 + 28h - 7h²       - 0

Arrange in descending powers of h

-7h² + 28h + 84 = 0

Divide through by -7

 h² - 4h - 12 = 0

Factor:

(h - 6)(h + 2) = 0

h - 6 = 0;   h + 2 = 0
    h = 6;       h = -2

To find the value of r corresponding to h = 6, substitute
in

  10 - 6h + h² - 2r = 0
10 - 6(6) + 6² - 2r = 0   
  10 - 36 + 36 - 2r = 0  
            10 - 2r = 5    
                -2r = -10
                  r = 5

So one possibility is the amaller circle 

(x - h)² + (y - r)² = r²

which becomes

(x - 6)² + (y - 5)² = 5²

That's the standard form, which when multiplied out gives
the general form:



x² + y² - 12x - 10y + 36 = 0

---
To find the value of r corresponding to h = -2,  substitute
h = 2 in

  10 - 6h + h² - 2r = 0
10 - 6(-2) + (-2)² - 2r = 0   
   10 + 12 + 4 - 2r = 0  
            26 - 2r =     
                -2r = -26
                  r = 13

And the other possibility is the larger circle which is

(x - h)² + (y - r)² = r²
(x + 2)² + (y - 13)² = 13²

That's the standard form, which when multiplied out gives
the general form:

x² + y² + 4x - 25y + 4 = 0

Two possible solutions.

Edwin